# Surfaces and other Peano Continua with no Generic Chains

### v1

@@ A.2. Chains on $S^{1}$
In this section we reduce to the results in the previous section to prove the analogous results for the circle.
@@
For each $z\in S^{1}$, let $\Phi_{z}(S^{1})$ be the closed subset of chains whose endpoints is $z$, and fix the clockwise $\rho_{z}$ rotation taking $z$ to $\pi(0)$. Let $\mathcal{R}\colon\Phi(S^{1})\to\Phi_{\pi(0)}(S^{1})$ be the continuous map:
@@ Appendix B The Sierpiński carpet
This appendix is devoted to proving that the action $\operatorname{Homeo}(S)\curvearrowright\Phi(S)$ is minimal, for the Sierpiński carpet $S$, and thus concluding that $\operatorname{Homeo}(S)$ does not have the generic point property and its universal minimal flow is not metrizable. Since, as we will see, the Sierpiński carpet is not homogeneous, and thus not strongly locally homogeneous, the minimality of $\operatorname{Homeo}(S)\curvearrowright\Phi(S)$ does not follow from Theorem 7.7. We do not settle whether this action is turbulent.
@@
As we can see, $S$ satisfies condition 1.2 of Theorem 1.2.
@@
The Sierpiński carpet has two types of points: rational and irrational. The collection of rational points is given by the union of all the boundaries of the maximal connected components of the complements of $S$ in the plane or, equivalently and more abstractly, it is the union of all non-trivial non locally separating proper subcontinua of $S$ (see [MR0200910]*§8). There are countably many maximal such subcontinua, called rational boundary curves, they are all simple closed curves, and their union is dense and co-dense in the space. Since it is the union of countably many non locally separating subcontinua, the set of rational points is itself non locally separating, by [MOT]*§2.2. The set of irrational points is the complement. A subcontinuum of $S$ is said to be rational/irrational if all of its points are rational/irrational.
@@
Let $\mathcal{W}$ be a brick partition of $S$ and $W\in\mathcal{W}$.
@@
(1). If $W=W^{\prime}$, take any irrational point in $W$. Otherwise, suppose towards contradiction that each point in $C\coloneqq\operatorname{Cl}(W)\cap\operatorname{Cl}(W^{\prime})\cap(W\vee W^{\prime})$ is rational in $S$ and thus also in $\operatorname{Cl}(W\cup W^{\prime})$, by [MR0749904]*Lemma 2. Since it is the countable union of non locally separating subcontinua, $C$ is non locally separating in $\operatorname{Cl}(W\cup W^{\prime})$. But $(W\vee W^{\prime})\setminus C=W\cup W^{\prime}$, which is disconnected, a contradiction.
(2). If $x,y$ are in $W$ and are irrational, then there is an irrational arc connecting $x,y$ because the set of rational points is non locally separating. By following the strategy laid out in [nadler2017continuum]*8.32(a), we then obtain the general statement.
@@
A basis of open sets in $C(S)$ is given by sets of the form $O(\mathcal{W}_{0})$, for an $E$-connected subset $\mathcal{W}_{0}\subseteq\mathcal{W}$ of some brick partition $\mathcal{W}$ on $S$. Let $w$ be a reduced walk on $\mathcal{W}$, which is confined to and visits all vertices of $\mathcal{W}_{0}$.
By (Upgrading Walks to Paths). (Theorem 6.2) and (Path Doubling). (Proposition 6.1), there are a brick partition $\mathcal{V}\preceq\mathcal{W}$ and two paths $v_{0},v_{1}$ on $\mathcal{V}$ which are disjoint except for their endpoints and such that $\rho^{\mathcal{V}}_{\mathcal{W}}\cdot v_{i}\sqsubseteq w$, for $i<2$. Let $v$ be the a circular path which follows $v_{0}$ forward and then $v_{1}$ backwards. For each $i<\ell$, let $x_{i}$ be an irrational point in $\operatorname{Cl}(v(i))\cap\operatorname{Cl}(v(i+1\text{ mod}\,\ell))\cap(v(i)\vee v(i+1\text{ mod}\,\ell))$, which exists by Proposition B.2.(1) . By point (2) of the same proposition, there are irrational arcs $\gamma_{i}$ in $v(i)\cup\{x_{i},x_{i+1\text{ mod}\,\ell}\}$ connecting $x_{i}$ to $x_{i+1\text{ mod}\,\ell}$. Let $\gamma$ be the union of such arcs. It is an irrational simple closed curve which meets each $v(i)$, for $i<\operatorname{lh}(v)$, and is contained in $\bigvee_{i<\operatorname{lh}(v)}v(i)$. Since $\rho^{\mathcal{V}}_{\mathcal{W}}\cdot v\sqsubseteq w$, it follows that $\gamma$ meets each $W\in\mathcal{W}_{0}$ and is contained in $\bigvee\mathcal{W}_{0}$, that is, $\gamma\in O(\mathcal{W}_{0})$. ∎
@@
Remark B.5.
Since $S$ is planar, each simple closed curve $\gamma\subseteq S$ separates $S$ into two components $W_{0},W_{1}$ (one of which could be empty). As noted in [whyburn]*§2, if every rational maximal subcontinuum of $S$ that is contained in $W_{0}\cup\gamma$ is either disjoint from $\gamma$ or coincides with $\gamma$, then $W_{0}\cup\gamma$ is an $S$-curve (and similarly for $W_{1}\cup\gamma$), and hence it is homeomorphic to $S$. This is the case for instance if $\gamma$ is irrational. A fact that we will often use implicitly is that a point in $W_{0}$ is irrational in $W_{0}\cup\gamma$ if and only if it is irrational in $S$.
@@
    $\theta,\zeta\subseteq U_{0}$ and $\eta\subseteq U_{1}$,
@@
Now, let $x\in S\setminus\operatorname{Cl}(W^{\prime}_{0})$ be given. Notice that $S\setminus W^{\prime}_{0}$ is an $S$-curve and that $\gamma,\gamma^{\prime}_{0}$ are rational boundary curves in it. By Lemma 6.8, there are a brick partition $\mathcal{V}^{\prime}$ of $S\setminus W^{\prime}_{0}$ and two disjoint spaced circular paths $c,c^{\prime}_{0}$ on $\mathcal{V}^{\prime}$ which are $\gamma$-compatible and $\gamma^{\prime}_{0}$-compatible, respectively. We can moreover suppose that $\mathcal{V}^{\prime}$ is fine enough so that:
  1. (1)
@@
which is possible by (Path Doubling). (Proposition 6.1.) Since $c$ is $\gamma$-compatible, $\lvert\gamma\cap c(i)\rvert\geq 2$, for each $i<\operatorname{lh}(c)$, and the same is true for $c^{\prime}_{0}$ and $\gamma^{\prime}_{0}$.
Using Proposition B.2.(2), we find an arc $\eta\subseteq c(i)$ whose endpoints belong to $\gamma$ and whose other points are irrational. In particular, $\eta\subseteq U_{1}$.
By Proposition B.2.(3), there are two disjoint arcs $\zeta_{0},\zeta_{1}$ in $\bigvee_{i\in\operatorname{lh}(P^{\prime})}P^{\prime}(i)$ which are irrational (in $S\setminus W^{\prime}_{0}$) except at their endpoint and connect $\gamma^{\prime}_{0}$ to $\gamma$. Now, $\gamma^{\prime}_{0}$ is split into two arcs by the endpoints of $\zeta_{0},\zeta_{1}$. One of these arcs, call it $\zeta_{2}$, is such that $\zeta\coloneqq\zeta_{0}\cup\zeta_{2}\cup\zeta_{1}$ separates $S$ into two components, one of which contains $W^{\prime}_{0}$ and is contained in $W^{\prime}_{0}\cup\bigvee_{i\in\operatorname{lh}(P^{\prime})}P^{\prime}(i)\subseteq U_{0}$. Call this component $V^{\prime}_{0}$. Then $W_{0}\subseteq V^{\prime}_{0}$, while $x$ and $U_{1}$ are in $S\setminus V^{\prime}_{0}$.
By the same reasoning as above, there are two disjoint arcs $\theta_{0},\theta$ in $V^{\prime}_{0}\cup\gamma$ which are irrational (in $S\setminus W_{0}$) except at their endpoint and connect $\gamma_{0}$ to $\gamma$. Now, $\gamma_{0}$ is split into two arcs by the endpoints of $\theta_{0},\theta_{1}$. One of these arcs, call it $\theta_{2}$, is such that $\theta\coloneqq\theta_{0}\cup\theta_{2}\cup\theta_{1}$ separates $S$ into two components, one of which contains $W_{0}$ and is contained in $V^{\prime}_{0}$. Call this component $V_{0}$.
Then $\theta\subseteq V^{\prime}_{0}$, while $x$ and $U_{1}$ are in $S\setminus V^{\prime}_{0}$, so point 2 is satisfied. On the other hand, $W_{0}\subseteq V_{0}$ and $\zeta\subseteq S\setminus V_{0}$, so also point 3 is taken care of. ∎
@@
Since $\operatorname{Homeo}(S)\curvearrowright S$ is minimal, it suffices to verify the conditions of Proposition 7.1. Let $\mathcal{C}\in\Phi(X)$ with root $x_{0}$ be given, together with open connected sets $U_{0},U_{1},V$ such that $x_{0}\in U_{0}$ and $U_{0},U_{1}\subseteq V$. In case $U_{0}\cap U_{1}\not=\emptyset$, by minimality of $\operatorname{Homeo}(S)\curvearrowright S$ we can assume that $x_{0}\in U_{0}\cap U_{1}$, so the conclusion follows. We henceforth consider the case $U_{0}\cap U_{1}=\emptyset$.
@@
    $\theta,\zeta\subseteq U_{0}$ and $\eta\subseteq U_{1}$,
@@
[…]
[scale=1]img/sierpi
Figure 3. The construction in the proof of Theorem B.7.
@@
The homeomorphisms $h_{0},h_{1},h_{2,3}$ all agree on their common domains so they give rise to a homeomorphism $h\colon\Sigma\to\Sigma$ that fixes set-wise $\gamma$. We can use one last time Theorem B.4 on $\tilde{h}\coloneqq h\upharpoonright\gamma$ to obtain a homeomorphism $h^{\prime}\colon\Sigma^{\prime}\to\Sigma^{\prime}$ which agrees with $h$ on $\gamma$. Their union $g$ is a homeomorphism of $S$ such that $g(x_{0})\in U_{0}$, $g(x_{1})\in U_{1}$ and $g[Z]=Z\subseteq V$, so in particular $g[C]\subseteq V$, as desired. ∎
Therefore we conclude that:
@@
$\operatorname{Homeo}(S)$, for the Sierpiński carpet $S$, does not have the generic point property. In particular, its universal minimal flow is not metrizable.

### v3

@@ A.2. Chains on $S^{1}$
In this section we reduce to the results in the previous section to prove analogous statements for the circle.
@@
For each $z\in S^{1}$, let $\Phi_{z}(S^{1})$ be the closed subset of chains whose endpoint is $z$, and fix the clockwise $\rho_{z}$ rotation taking $z$ to $\pi(0)$. Let $\mathcal{R}\colon\Phi(S^{1})\to\Phi_{\pi(0)}(S^{1})$ be the continuous map:
@@ Appendix B The Sierpiński carpet
This appendix is devoted to proving that the action $\operatorname{Homeo}(S)\curvearrowright\Phi(S)$ is minimal, for the Sierpiński carpet $S$, and thus concluding that $\operatorname{Homeo}(S)$ does not have the generic point property and its universal minimal flow is not metrizable. Since, as we will see, the Sierpiński carpet is not homogeneous, and thus not strongly locally homogeneous, the minimality of $\operatorname{Homeo}(S)\curvearrowright\Phi(S)$ does not follow from Theorem 7.7. We do not settle whether this action is generically turbulent.
@@
It is immediate that $S$ satisfies condition (2) of Theorem 1.2, and it is not hard to see that it also satisfies condition (3).
@@
The Sierpiński carpet has two types of points: rational and irrational. The collection of rational points is given by the union of all the boundaries of the maximal connected components of the complements of $S$ in the plane or, equivalently and more abstractly, it is the union of all non-trivial non-locally-separating proper subcontinua of $S$ (see [MR0200910]*§8). There are countably many maximal such subcontinua, called rational boundary curves, they are all simple closed curves, and their union is dense and co-dense in the space. Since it is the union of countably many non-locally-separating subcontinua, the set of rational points is a non-locally-separating $F_{\sigma}$ set, by [MOT]*Theorem 2.2. The set of irrational points is the complement. A subcontinuum of $S$ is said to be rational/irrational if all of its points are rational/irrational.
@@
Let $\mathcal{W}$ be a brick partition of $S$ and $W,W^{\prime}\in\mathcal{W}$.
@@
(1) and (2) follow directly from 4.3.(4) and 4.2.(3), respectively.
@@
[…]
By (Upgrading Walks to Paths). (Theorem 6.2) and (Path Doubling). (Proposition 6.1), there are a brick partition $\mathcal{V}\preceq\mathcal{W}$ and two paths $v_{0},v_{1}$ on $\mathcal{V}$ which are disjoint except for their endpoints and such that $\rho^{\mathcal{V}}_{\mathcal{W}}\cdot v_{i}\sqsubseteq w$, for $i<2$. Let $v$ be the circular path which follows $v_{0}$ forward and then $v_{1}$ backwards. For each $i<\ell$, let $x_{i}$ be an irrational point in $\operatorname{Cl}(v(i))\cap\operatorname{Cl}(v(i+1\text{ mod}\,\ell))\cap(v(i)\vee v(i+1\text{ mod}\,\ell))$, which exists by Proposition B.2.(1). By point (2) of the same proposition, there are irrational arcs $\gamma_{i}$ in $v(i)\cup\{x_{i},x_{i+1\text{ mod}\,\ell}\}$ connecting $x_{i}$ to $x_{i+1\text{ mod}\,\ell}$. Let $\gamma$ be the union of such arcs. It is an irrational simple closed curve which meets each $v(i)$, for $i<\operatorname{lh}(v)$, and is contained in $\bigvee_{i<\operatorname{lh}(v)}v(i)$. Since $\rho^{\mathcal{V}}_{\mathcal{W}}\cdot v\sqsubseteq w$, it follows that $\gamma$ meets each $W\in\mathcal{W}_{0}$ and is contained in $\bigvee\mathcal{W}_{0}$, that is, $\gamma\in O(\mathcal{W}_{0})$. ∎
@@
Let $\gamma$ be a simple closed curve in $S$, and let $W$ be a connected component of $S\setminus\gamma$. As noted in [whyburn]*§2, if every rational maximal subcontinuum of $S$ that is contained in $W\cup\gamma$ is either disjoint from $\gamma$ or coincides with $\gamma$, then $W\cup\gamma$ is an $S$-curve, and hence it is homeomorphic to $S$. This is the case for instance if $\gamma$ is irrational and $W$ is any one of the two components of $S\setminus\gamma$. A fact that we will often use implicitly is that a point in $W$ is irrational in $W\cup\gamma$ if and only if it is irrational in $S$.
@@
    $\theta,\zeta\subseteq U_{0}$, one of the connected components of $S\setminus\theta$ is contained in $U_{0}$, and the same is true of one of the connected components of $S\setminus\zeta$,
  2. (2)
    $\eta\subseteq U_{1}$, and one of the connected components of $S\setminus\eta$ is contained in $U_{1}$,
@@
Now, let $x\in S\setminus\operatorname{Cl}(W^{\prime}_{0})$ be given. Notice that $\Sigma\coloneqq S\setminus W^{\prime}_{0}$ is an $S$-curve and that $\gamma,\gamma^{\prime}_{0}$ are rational boundary curves in it. By Lemma 6.8, there are a brick partition $\mathcal{V}^{\prime}$ of $\Sigma$ and two disjoint spaced circular paths $c,c^{\prime}_{0}$ on $\mathcal{V}^{\prime}$ which are $\gamma$-compatible and $\gamma^{\prime}_{0}$-compatible, respectively. We can moreover suppose that $\mathcal{V}^{\prime}$ is fine enough so that:
  1. (i)
@@
which is possible by (Path Doubling). (Proposition 6.1). Since $c$ is $\gamma$-compatible, $\lvert\gamma\cap c(i)\rvert\geq 2$, for each $i<\operatorname{lh}(c)$, and the same is true for $c^{\prime}_{0}$ and $\gamma^{\prime}_{0}$.
Using Proposition B.2.(2), we find an arc $\eta\subseteq c(i)$ whose endpoints belong to $\gamma$ and whose other points are irrational. In particular, $\eta\subseteq U_{1}$, and since $\gamma\cap c(i)$ is connected, one the connected components of $S\setminus(\gamma\cup\eta)$ is contained in $U_{1}$.
By Proposition B.2.(3), there are two disjoint arcs $\zeta_{0},\zeta_{1}$ in $\bigvee_{i\in\operatorname{lh}(P^{\prime})}P^{\prime}(i)$ which are irrational (in $\Sigma$) except at their endpoint and connect $\gamma^{\prime}_{0}$ to $\gamma$. Now, $\gamma^{\prime}_{0}$ is split into two arcs by the endpoints of $\zeta_{0},\zeta_{1}$. One of these arcs, call it $\zeta_{2}$, is such that $\zeta\coloneqq\zeta_{0}\cup\zeta_{2}\cup\zeta_{1}$ separates $S$ into two components, one of which contains $W^{\prime}_{0}$ and is contained in $W^{\prime}_{0}\cup\bigvee_{i\in\operatorname{lh}(P^{\prime})}P^{\prime}(i)\subseteq U_{0}$. Call this component $V^{\prime}_{0}$. Then $W_{0}\subseteq V^{\prime}_{0}$, while $x$ and $U_{1}$ are in $S\setminus V^{\prime}_{0}$.
By the same reasoning as above, there are two disjoint arcs $\theta_{0},\theta_{1}$ in $V^{\prime}_{0}\cup\gamma$ which are irrational (in $S\setminus W_{0}$) except at their endpoint and connect $\gamma_{0}$ to $\gamma$. Now, $\gamma_{0}$ is split into two arcs by the endpoints of $\theta_{0},\theta_{1}$. One of these arcs, call it $\theta_{2}$, is such that $\theta\coloneqq\theta_{0}\cup\theta_{2}\cup\theta_{1}$ separates $S$ into two components, one of which contains $W_{0}$ and is contained in $V^{\prime}_{0}$. Call this component $V_{0}$.
Then $\theta\subseteq V^{\prime}_{0}$, while $x$ and $U_{1}$ are in $S\setminus V^{\prime}_{0}$, so point (3) is satisfied. On the other hand, $W_{0}\subseteq V_{0}$ and $\zeta\subseteq S\setminus V_{0}$, so also point (4) is taken care of. ∎
@@
Since $\operatorname{Homeo}(S)\curvearrowright S$ is minimal, it suffices to verify the conditions of Proposition 7.1. Let $\mathcal{C}\in\Phi(X)$ with root $x_{0}$ be given, together with nonempty open connected sets $U_{0},U_{1},V$ such that $U_{0},U_{1}\subseteq V$. In case $U_{0}\cap U_{1}\not=\emptyset$, by minimality of $\operatorname{Homeo}(S)\curvearrowright S$ we can assume that $x_{0}\in U_{0}\cap U_{1}$, so the conclusion follows. We henceforth consider the case $U_{0}\cap U_{1}=\emptyset$.
@@
    $\theta,\zeta\subseteq U_{0}$, one of the connected components of $S\setminus\theta$ is contained in $U_{0}$, and the same is true of one of the connected components of $S\setminus\zeta$,
  2. (2)
    $\eta\subseteq U_{1}$, and one of the connected components of $S\setminus\eta$ is contained in $U_{1}$,
@@
[…]
[scale=1]img/sierpi
Figure 3. The construction in the proof of Theorem B.6.
@@
The homeomorphisms $h_{0},h_{1},h_{2,3}$ all agree on their common domains, so they give rise to a homeomorphism $h\colon\Sigma\to\Sigma$ that fixes set-wise $\gamma$. We can use one last time Theorem B.4 on $\tilde{h}\coloneqq h\upharpoonright\gamma$ to obtain a homeomorphism $h^{\prime}\colon\Sigma^{\prime}\to\Sigma^{\prime}$ which agrees with $h$ on $\gamma$. Their union $g$ is a homeomorphism of $S$ such that $g(x_{0})\in U_{0}$, $g(x_{1})\in U_{1}$ and $g[Z]=Z\subseteq V$, so in particular $g[C]\subseteq V$, as desired. ∎
Therefore, we conclude that:
@@
$\operatorname{Homeo}(S)$, for the Sierpiński carpet $S$, does not have the generic point property. In particular, its universal minimal flow is not metrizable and has no comeager orbit.